# AverageTime: Enhance Long-Term Time Series Forecasting with Simple Averaging

### v3

@@ 4 Experiments
The experimental section is organized as follows: we begin by introducing the benchmark datasets we utilized, which are widely used in long-term time series forecasting. We then present the baselines used for comparison. Next, we describe the experimental setup and provide the main experimental results for AverageTime and LightAverageTime. Following this, we discuss the impact of hyperparameters on LightAverageTime as well as the effect of lookback windows on model performance. Afterward, we conduct a complexity analysis to assess the computational efficiency of our model. Finally, we present ablation studies to investigate the impact of the average framework on the model’s performance. All experiments in this section were conducted on a single NVIDIA RTX 4090D GPU.
@@ 4.1 Datasets
In this section, we first introduce the datasets used in our method, which include eight classic datasets for long-term time series forecasting. Their basic attributes are shown in Table 1:
@@
Data ETTh1 ETTh2 ETTm1 ETTm2 Weather Solar ECL Traffic
Channels 7 7 7 7 21 137 321 862
Frequency 1 hour 1 hour 15 mins 15 mins 10 mins 10 mins 1 hour 1 hour
Steps 17,420 17,420 69,680 69,680 52,696 52,179 26,304 17,544
Normally, the ETT datasets have fewer channels, making them easier to predict even with simple parameter-independent MLP layers. In contrast, for datasets with more channels, the average framework tends to be more effective due to its ability to effectively capture channel dependencies.
@@ 4.4 Main results
In this section, we compare the primary results of AverageTime and LightAverageTime against the baseline models mentioned in Section 4.2 for long-term time series forecasting. Besides our proposed methods, the experimental results for other approaches are sourced from iTransformer (Liu et al., 2024) and TimeXer (Wang et al., 2024b). Specifically, for the Solar dataset, we implemented the TimeXer and CycleNet methods on a 4090D GPU ourselves and adjusted the hyperparameters to achieve better results, as their original papers did not provide these results. The results are then averaged over different prediction lengths. We bold the best and underline the second-best evaluation result for each dataset. The detailed comparison is presented in Table 2.
@@
Datasets Avg.Time LAvg.Time TimeXer CycleNet iTrans. RLinear PatchTST TimesNet
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 $\underline{0.438}$ 0.431 0.437 0.431 0.437 0.437 0.457 0.441 0.454 0.448 0.446 $\underline{0.434}$ 0.469 0.455 0.458 0.450
ETTh2 $\underline{0.372}$ $\underline{0.398}$ 0.377 0.401 0.368 0.396 0.388 0.409 0.383 0.407 0.374 0.399 0.387 0.407 0.414 0.427
ETTm1 $\underline{0.381}$ 0.391 0.383 $\underline{0.393}$ 0.382 0.397 0.379 0.396 0.410 0.410 0.414 0.408 0.390 0.400 0.398 0.406
ETTm2 $\underline{0.274}$ $\underline{0.318}$ 0.275 0.319 $\underline{0.274}$ 0.322 0.266 0.314 0.288 0.332 0.286 0.327 0.281 0.336 0.291 0.333
Weather 0.242 $\underline{0.271}$ 0.241 0.269 0.241 $\underline{0.271}$ 0.243 0.271 0.258 0.278 0.272 0.291 0.259 0.281 0.259 0.287
Solar $\underline{0.229}$ 0.271 0.237 0.273 0.233 0.281 0.210 $\underline{0.263}$ 0.233 0.262 0.369 0.357 0.254 0.289 0.301 0.319
ECL 0.174 $\underline{0.270}$ 0.176 0.272 $\underline{0.171}$ $\underline{0.270}$ 0.168 0.259 0.178 $\underline{0.270}$ 0.219 0.298 0.216 0.304 0.193 0.295
Traffic 0.424 0.279 $\underline{0.426}$ $\underline{0.281}$ 0.466 0.286 0.472 0.301 0.428 0.282 0.627 0.378 0.474 0.304 0.620 0.336
Average 0.317 0.329 $\underline{0.319}$ $\underline{0.330}$ 0.322 0.332 0.323 0.332 0.329 0.336 0.376 0.361 0.342 0.355 0.367 0.357
@@
AverageTime achieved the best model performance among all compared models, with a 1.55% reduction in MSE and a 0.90% reduction in MAE compared with the second-best model, TimeXer, demonstrating the excellent performance of our model. The lightweight version, LightAverageTime with a cluster block, also outperforms TimeXer. Notably, the performance of AverageTime and TimeXer is nearly identical across the first seven datasets, with identical MSE performance and a 0.88% improvement in MAE. However, the mixed-channel embedding layer significantly improves our model’s performance on the Traffic dataset, indicating that AverageTime is more robust than TimeXer.
For datasets with a smaller number of channels, linear models like RLinear often exhibit sufficient predictive capability, as evidenced by the model’s superior performance on the ETT scenario. However, linear models typically fail in datasets with more channels. Moreover, our average framework is more effective for these scenarios. Notably, through mixed-channel embedding within the averaging framework, our model matched or even surpassed the performance of state-of-the-art Transformer architectures on Weather, ECL, and Traffic datasets, and the overall performance significantly outperforms Linear models like RLinear. This indicates that the average framework is simple yet of great importance in long-term time series forecasting.
@@ 4.5 Detail for LightAverageTime
As described in section 3.2, LightAverageLinear is highly dependent on the hyperparameter we define to filter out channels with low correlation to all other channels. For a fair comparison, we set the threshold value to 0.8 in the main results. The performance of LightAverageLinear can vary significantly depending on the selected threshold. When the threshold value approaches 1, the model behaves more like a parameter-independent method, whereas when the threshold is close to 0, it functions as a parameter-dependent method. Thus, adjusting the threshold can significantly affect the model’s predictions. The efficiency and performance of the model for different threshold values are illustrated in Figure 3:
(a) Different thresohold on model performance and efficiency on weather dataset.
(b) Different threshold on model performance and efficiency on electricity dataset.
Figure 3: Different threshold on model performance and efficiency on different datasets.
@@ 4.6 Complexity Analysis
(a) ETTh1 efficiency.
(b) Weather efficiency.
Figure 4: Comparison of Model Complexities: We compare the complexities of different models. AverageTime and LightAverageTime demonstrate efficiency comparable to that of linear models, significantly outperforming Transformer-based models with approximately half the memory usage and 70% of the training time compared to the second-best model, TimeXer.
@@
4.7 Lookback windows experiment
Models with different lookback windows show varying performance. Some models that perform well with a lookback window of 96 may underperform with other lookback windows. To validate that the performance of AverageTime can be extended to different lookback windows, we conduct experiments on four datasets—ETTh1, ETTm1, weather, and traffic, with channel sizes ranging from 7 to 862. Hyperparameters are kept consistent with those used for a lookback window of 96 to simplify the experiment. The average results for different prediction lengths across the datasets are shown in Figure 5:
Figure 5: The model’s performance varies with changes in the lookback window.
It can be noted that, compared to the model with a lookback window of 96, the performance of AverageTime is consistently improved with longer lookback windows, which demonstrates that the average framework is effective across different lookback windows. The best performance is typically achieved when the prediction length equals 336. Note that the hyperparameters are consistent with those used for the model with a lookback window of 96. Since the model’s performance can be influenced by hyperparameters, carefully adjusting them can further enhance the model’s performance.
@@ 4.8 Ablation Analysis
In this section, we conduct an ablation study to evaluate the effectiveness of the averaging framework. Specifically, in the comparison model, we remove all channel embedding layers, leaving only simple MLP layers for prediction while keeping the hyperparameters consistent. The results are presented in Table 3.
Table 3: Ablation study on average framework.
Datasets MLP +Average Improv.(%)
Metrics MSE MAE MSE MAE MSE MAE
ETTh1 0.440 0.433 0.438 0.431 0.45 0.46
ETTh2 0.390 0.409 0.372 0.398 4.62 2.69
ETTm1 0.386 0.395 0.381 0.391 1.30 1.01
ETTm2 0.276 0.320 0.274 0.318 0.72 0.63
Weather 0.247 0.274 0.242 0.271 2.07 1.09
Solar 0.303 0.308 0.229 0.271 24.42 12.01
ECL 0.195 0.284 0.174 0.270 10.77 4.93
Traffic 0.517 0.325 0.424 0.279 17.99 14.15
Average 0.344 0.344 0.317 0.329 6.98 4.36
It can be observed that the averaging framework significantly enhances the model’s performance, particularly in datasets with a large number of channels, such as Solar, ECL, and Traffic. However, its effectiveness diminishes in datasets with fewer channels, though it still provides some improvement. This suggests that the impact of the averaging operation is more pronounced in high-channel settings and plays a crucial role in performance enhancement.

### v4

@@ 4 Experiments
The experimental section is organized as follows: we begin by introducing the benchmark datasets we utilized, which are widely used in long-term time series forecasting. We then present the baselines used for comparison. Next, we describe the experimental setup and provide the main experimental results for AverageTime and LightAverageTime. We further conduct a detailed analysis of the information gain brought by the original and extracted time series of AverageTime to demonstrate the effectiveness of our method. Following this, we discuss how the FVMD-decomposed time series can be integrated into the AverageTime framework, and the impact of hyperparameters on LightAverageTime. Afterward, we conduct a complexity analysis to assess the computational efficiency of our model. All experiments in this section were conducted on a single NVIDIA RTX 4090D GPU.
@@ 4.1 Datasets
In this section, we first introduce the datasets used in our method, which include seven classic datasets for long-term time series forecasting. Their basic attributes are shown in Table 1:
@@
Data ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL Traffic
Channels 7 7 7 7 21 321 862
Frequency 1 hour 1 hour 15 mins 15 mins 10 mins 1 hour 1 hour
Steps 17,420 17,420 69,680 69,680 52,696 26,304 17,544
Normally, the ETT datasets have fewer channels, making them easier to predict even with only parameter-independent MLP layers. For datasets with more channels, the averaging framework becomes more effective because extracting information from more channels yields time series with richer information.
@@ 4.4 Main results
In this section, we compare the primary results of AverageTime and LightAverageTime against the baseline models mentioned in Section 4.2 for long-term time series forecasting. Besides our proposed methods, the experimental results for other approaches are sourced from iTransformer [9] and TimeXer [16]. We bold the best and underline the second-best evaluation result for each dataset. The detailed comparison is presented in Table 2.
@@
Avg.Time LAvg.Time TimeXer CycleNet iTrans. RLinear PatchTST TimesNet
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.378 $\underline{0.393}$ $\underline{0.376}$ 0.392 0.382 0.403 0.375 0.395 0.386 0.405 0.386 0.395 0.414 0.419 0.384 0.402
ETTh1 192 $\underline{0.434}$ 0.424 0.435 0.425 0.429 0.435 0.436 $\underline{0.428}$ 0.441 0.436 0.437 0.424 0.460 0.445 0.436 0.429
336 $\underline{0.475}$ $\underline{0.446}$ 0.474 0.445 0.468 0.448 0.496 0.455 0.487 0.458 0.479 $\underline{0.446}$ 0.501 0.466 0.491 0.469
720 $\underline{0.466}$ $\underline{0.461}$ 0.464 0.460 0.469 $\underline{0.461}$ 0.520 0.484 0.503 0.491 0.481 0.470 0.500 0.488 0.521 0.500
96 $\underline{0.290}$ $\underline{0.340}$ 0.292 0.343 0.286 0.338 0.298 0.344 0.297 0.349 0.288 0.338 0.302 0.348 0.340 0.374
ETTh2 192 $\underline{0.370}$ $\underline{0.390}$ 0.376 0.393 0.363 0.389 0.372 0.396 0.380 0.400 0.374 0.390 0.388 0.400 0.402 0.414
336 0.414 $\underline{0.426}$ 0.417 0.428 0.414 0.423 0.431 0.439 0.428 0.432 $\underline{0.415}$ $\underline{0.426}$ 0.426 0.433 0.452 0.452
720 $\underline{0.413}$ $\underline{0.435}$ 0.423 0.441 0.408 0.432 0.450 0.458 0.427 0.445 0.420 0.440 0.431 0.446 0.462 0.468
96 0.314 0.353 $\underline{0.318}$ 0.358 $\underline{0.318}$ $\underline{0.356}$ 0.319 0.360 0.334 0.368 0.355 0.376 0.329 0.367 0.338 0.375
ETTm1 192 0.356 0.375 $\underline{0.360}$ $\underline{0.378}$ 0.362 0.383 0.360 0.381 0.387 0.391 0.391 0.392 0.367 0.385 0.374 0.387
336 $\underline{0.392}$ 0.398 $\underline{0.392}$ $\underline{0.400}$ 0.395 0.407 0.389 0.403 0.426 0.420 0.424 0.415 0.399 0.410 0.410 0.411
720 0.463 $\underline{0.438}$ 0.460 0.435 $\underline{0.452}$ 0.441 0.447 0.441 0.491 0.459 0.487 0.450 0.454 0.439 0.478 0.450
96 $\underline{0.170}$ $\underline{0.251}$ $\underline{0.170}$ 0.253 0.171 0.256 0.163 0.246 0.180 0.264 0.182 0.265 0.175 0.259 0.187 0.267
ETTm2 192 $\underline{0.235}$ $\underline{0.294}$ 0.236 0.295 0.237 0.299 0.229 0.290 0.250 0.309 0.246 0.304 0.241 0.302 0.249 0.309
336 $\underline{0.293}$ $\underline{0.332}$ 0.297 0.334 0.296 0.338 0.284 0.327 0.311 0.348 0.307 0.342 0.305 0.343 0.321 0.351
720 0.396 $\underline{0.393}$ 0.396 $\underline{0.393}$ $\underline{0.392}$ 0.394 0.389 0.391 0.412 0.407 0.407 0.398 0.402 0.440 0.408 0.403
96 0.156 0.203 0.156 0.203 $\underline{0.157}$ $\underline{0.205}$ 0.158 0.203 0.174 0.214 0.192 0.232 0.177 0.218 0.172 0.220
Weather 192 $\underline{0.205}$ 0.248 $\underline{0.205}$ 0.245 0.204 $\underline{0.247}$ 0.207 $\underline{0.247}$ 0.221 0.254 0.240 0.271 0.225 0.259 0.219 0.261
336 0.263 $\underline{0.289}$ 0.260 0.286 $\underline{0.261}$ 0.290 0.262 $\underline{0.289}$ 0.278 0.296 0.292 0.307 0.278 0.297 0.280 0.306
720 0.345 0.345 $\underline{0.343}$ $\underline{0.342}$ 0.340 0.341 0.344 0.344 0.358 0.349 0.364 0.353 0.354 0.348 0.365 0.359
96 0.142 $\underline{0.240}$ 0.144 0.242 $\underline{0.140}$ 0.242 0.136 0.229 0.148 $\underline{0.240}$ 0.201 0.281 0.195 0.285 0.168 0.272
ECL 192 0.161 0.256 0.163 0.258 $\underline{0.157}$ 0.256 0.152 0.244 0.162 $\underline{0.253}$ 0.201 0.283 0.199 0.289 0.184 0.289
336 0.177 0.275 0.181 0.278 $\underline{0.176}$ 0.275 0.170 0.264 0.178 $\underline{0.269}$ 0.215 0.298 0.215 0.305 0.198 0.300
720 0.217 0.310 0.217 0.310 0.211 $\underline{0.306}$ $\underline{0.212}$ 0.299 0.225 0.317 0.257 0.331 0.256 0.337 0.220 0.320
96 0.394 0.263 0.396 $\underline{0.265}$ 0.428 0.271 0.458 0.296 $\underline{0.395}$ 0.268 0.649 0.389 0.462 0.295 0.593 0.321
Traffic 192 0.415 0.274 0.418 $\underline{0.276}$ 0.448 0.282 0.457 0.294 $\underline{0.417}$ 0.276 0.601 0.366 0.466 0.296 0.617 0.336
336 0.428 0.280 $\underline{0.430}$ $\underline{0.281}$ 0.473 0.289 0.470 0.299 0.433 0.283 0.609 0.369 0.482 0.304 0.629 0.336
720 0.457 0.297 $\underline{0.460}$ $\underline{0.300}$ 0.516 0.307 0.502 0.314 0.467 0.302 0.647 0.387 0.514 0.322 0.640 0.350
Count 19 23 11 15 17 10 11 12 2 3 1 4 0 0 0 0
Average 0.329 0.337 0.331 0.338 0.334 0.340 0.339 0.341 0.343 0.347 0.377 0.362 0.354 0.355 0.376 0.362
@@
AverageTime achieved the best model performance among all compared models, with a 1.50% reduction in MSE and a 0.88% reduction in MAE compared with the second-best model, TimeXer, demonstrating the excellent performance of our model. The lightweight version, LightAverageTime with a cluster block, also outperforms TimeXer. It is noteworthy that AverageTime achieved top-2 rankings in 42 out of all 56 metrics, significantly outperforming all other models. This demonstrates the strong stability of AverageTime, which maintains consistently robust forecasting performance across datasets with varying channel counts—whether on the ETT series with fewer channels, the medium-channel weather dataset, or the high-channel Traffic dataset.
We further compare AverageTime with the linear model RLinear. Compared to RLinear, AverageTime incorporates a module for extracting and fusing channel information. On the ETTh-series datasets, which have a longer temporal span and fewer channels, the performance gap between the two models is relatively small. This suggests that for such datasets, the channel module provides limited extractable information. In contrast, for data with a shorter temporal span and more channels, where inter-channel dependencies are more pronounced, the extraction of channel information significantly enhances the model’s forecasting performance.
We further compared the model’s prediction performance on the Ettm2 dataset, conducting a detailed comparison with three methods: TimeXer, PatchTST, and DLinear. The prediction results of different methods across varying forecast horizons are illustrated in the Figure 3:
Figure 3: Prediction performance comparison across models and forecast horizons. Rows (top to bottom): AverageTime, TimeXer, PatchTST, DLinear. Columns (left to right): Forecast horizons of 96, 192, 336, and 720 steps.
It can be observed that none of the four models achieve perfect prediction, with PatchTST and DLinear exhibiting larger prediction deviations. Between TimeXer and AverageTime, AverageTime demonstrates more accurate overall forecasting performance, capturing trends and details more effectively, and thus delivers more precise predictions than TimeXer. Furthermore, it is noted that state-of-the-art time series forecasting models still show considerable prediction bias across many datasets. Under this premise, we argue that providing richer informational input to the model holds higher priority than designing more complex model architectures. Therefore, we introduce series decomposition information to further enhance the model’s predictive capability. The decomposed sequence information is first fused with features extracted across channels and then integrated with the original series. The corresponding prediction results are presented in the Table 3:
Table 3: Comparison of forecasting performance averaged over four horizons on seven datasets.
FVMD+Avg.Time Avg.Time LAvg.Time TimeXer Improvement (%)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.434 0.428 0.438 0.431 0.437 0.431 0.437 0.437 +0.91 +0.70
ETTh2 0.369 0.397 0.372 0.398 0.377 0.401 0.368 0.396 +0.81 +0.25
ETTm1 0.380 0.391 0.381 0.391 0.383 0.393 0.382 0.397 +0.26 +0.00
ETTm2 0.273 0.317 0.274 0.318 0.275 0.319 0.274 0.322 +0.37 +0.31
Weather 0.241 0.269 0.242 0.271 0.241 0.269 0.241 0.271 +0.41 +0.74
Electricity 0.175 0.271 0.174 0.270 0.176 0.272 0.171 0.270 -0.58 -0.37
Traffic 0.423 0.278 0.424 0.279 0.426 0.281 0.466 0.287 +0.24 +0.36
Using four-decimal averaged results, AverageTime with FVMD achieves consistent improvements over Avg.Time on most datasets, with up to 0.84% MSE and 0.70% MAE reduction on ETTh1. Furthermore, the hyperparameters were adopted directly from the final AverageTime setup without any additional hyperparameter optimization; the reported performance gain is therefore reliable.
4.5 Evaluating the Information Gain of Original and Extracted Representations
To validate the effectiveness of our proposed method, we further conduct a detailed analysis of the information gained from the original and the extracted sequences. To this end, we design two experiments. First, we verify that the sequences extracted through channel processing enable predictions with richer information, thereby enhancing the model’s performance. We evaluate the model on the Weather and Electricity datasets and examine how its performance changes as a function of the number of MLP and Transformer layers in the channel information extraction module. The results are presented in Figure 4.
Figure 4: Effects of Layer Depth on Channel Information Extraction Performance (Top: Transformer layers on Weather & ECL datasets; Bottom: MLP layers on the same datasets)
It can be observed that the overall performance of the model incorporating the MLPs and Transformers channel information extraction layer is superior. Compared with the model without channel information extraction module, a significant improvement is achieved, indicating that extracting informative representations from the feature dimension can substantially enhance model effectiveness. Furthermore, as the number of channel embedding layers increases, more complex channel information can be captured, leading to a further improvement in model performance.
In addition, if only the sequences extracted with channel information are utilized for forecasting, a considerable amount of important information contained in the original sequence may be lost. To examine whether information loss occurs during the channel extraction process, we conducted an experiment that uses only the extracted sequence for prediction, excluding the original sequence information. This setup also allows us to verify whether incorporating the original sequence can mitigate such information loss, the results are shown in Table 4:
Table 4: Prediction performance with or without original sequence.
Weather Electricity
Horizon MSE (w/o) MSE (w/) MAE (w/o) MAE (w/) MSE (w/o) MSE (w/) MAE (w/o) MAE (w/)
96 0.171 0.156 0.220 0.203 0.145 0.142 0.247 0.240
192 0.217 0.203 0.259 0.248 0.165 0.161 0.262 0.256
336 0.273 0.205 0.299 0.289 0.183 0.177 0.281 0.275
720 0.353 0.248 0.352 0.345 0.224 0.217 0.318 0.310
It can be observed that, on the two representative datasets, Weather and Electricity, the models incorporating the original sequences outperform those without them across all prediction length. This indicates that during the channel extraction stage, part of the information contained in the original sequences is indeed lost. Therefore, integrating the original sequences into the model architecture and blending their information with the extracted sequences is essential to achieve better overall sequence modeling performance.
@@ 4.6 Detail for LightAverageTime
As described in section 3.2, LightAverageTime is highly dependent on the hyperparameter $T$ we define to filter out channels with low correlation to all other channels, where $T$ is the similarity threshold above which a connection (edge) is formed between channels, ultimately governing the structure of the community cluster matrix. To avoid potential reliability issues caused by excessive hyperparameter tuning, we uniformly set $T=0.8$ in all experimental configurations. The performance of LightAverageTime can vary significantly depending on the selected threshold. When the threshold value $T$ approaches 1, the model behaves more like a parameter-independent method, whereas when the $T$ is close to 0, it functions as a parameter-dependent method. Thus, adjusting the threshold can significantly affect the model’s predictions. The efficiency and performance of the model for different threshold values are illustrated in Figure 5:
Figure 5: Different thresholds on model performance and efficiency on different datasets, LightAverageTime achieves the same performance as TimeXer and iTransformer with better efficiency.
@@ 4.7 Complexity Analysis
Figure 6: Comparison of model complexities. AverageTime and LightAverageTime demonstrate efficiency comparable to that of linear models, significantly outperforming Transformer-based models with approximately half the memory usage and 70% of the training time compared to the second-best model, TimeXer.
@@ 4.8 Ablation Analysis
In this section, we conduct a detailed ablation study to examine the contributions of different components of the model and to explore how their interactions affect overall performance. We evaluate its three key components: the Revin operation, the parameter independent prediction mechanism, and the averaging operation after channel information extraction. Experiments are carried out on three datasets—ETTm1, Weather, and Electricity—and the results are summarized in the Table 5.
Table 5: Ablation study results of AverageTime on ETTm1, Weather, and Electricity datasets. The three-digit code indicates the presence (1) or absence (0) of the following components in order: Channel Independent (C), Revin (R), and Average (A). The best MSE and MAE results for each dataset are highlighted in bold. The bottom row shows the relative improvement of the best model (111) over the second-best.
Components ETTm1 Weather Electricity
MSE MAE MSE MAE MSE MAE
000 (None) 0.3030 0.3396 0.2918 0.3075 0.2191 0.3061
100 (C) 0.2973 0.3348 0.2632 0.2905 0.1946 0.2894
010 (R) 0.3022 0.3403 0.2940 0.3070 0.2108 0.2945
001 (A) 0.3004 0.3420 0.2819 0.3122 0.1976 0.2966
110 (C+R) 0.2957 0.3342 0.2660 0.2903 0.1876 0.2800
101 (C+A) 0.2995 0.3407 0.2593 0.2947 0.1942 0.2949
011 (R+A) 0.2980 0.3372 0.2771 0.2996 0.1891 0.2840
111 (All) 0.2934 0.3327 0.2628 0.2893 0.1811 0.2780
Improv. +0.78% +0.45% -1.35% +0.35% +3.47% +0.71%
The results indicate that almost every individual component contributes to performance improvement, and the combination of different modules further enhances model efficacy. In 5 out of the 6 metrics, the best predictive performance is achieved when all modules are employed; the remaining top result is attained by the model using only the channel-independent and averaging strategies, which further validates the broad effectiveness of the averaging operation. We also compared the best model (using all modules) against the second-best results from all other configurations. A significant improvement of up to 3.47% in MSE is observed on the Electricity dataset, demonstrating that all three architectural components are crucial to the model’s performance.